OpenStudy (anonymous):

Evaluate \[\int\limits_{C}^{} xydy-y ^{2}dx\] where C is the closed curve consisting of the top half of the circle \[x^{2}+y^{2}=4\] and the x-axis segment \[-2 \le x \le 2.\]

OpenStudy (anonymous):

Without using Green's theorem, you can parameterize the curve as \[C=C_i=\begin{cases} \begin{cases} x=2\cos t\\ y=2\sin t\\ 0\le t\le \pi \end{cases}&\text{for }i=1 \\\\ \begin{cases} x=t\\ y=0\\ -2\le t\le2 \end{cases}&\text{for }i=2 \end{cases}\] Then evaluate the integral as \[\int_C=\int_{C_1}+\int_{C_2}\]

OpenStudy (anonymous):

If you have the luxury of using Green's theorem, you need to find the right partial derivatives. \[\int_C M(x,y)~dx+N(x,y)~dy=\int\int_D \left(\frac{\partial N}{\partial x}-\frac{\partial M}{\partial y}\right)~dA\] where \(D\) would be the region bounded by the curve \(C\).

OpenStudy (anonymous):

Thanks!